I am trying to understand calculating distance using the Pythagorean Theorem. An example of a problem is find the distance between (-1,3),(11,-2). I cam up with 13.6 but it doesn't feel right. PLEASE HELP!

Answer by richwmiller(17219)

this is a classic 5, 12, 13 right triangle
one leg is 3+2=5
the other leg is 1+11=12
You should immediately recognize this as a 5,12,13 right triangle
but if you didn't
25+144=169
5^2+12^2=13^2